# Standing-Wave Dynamics in Low-Frequency Breathing of a Turbulent Separation Bubble

### v2

@@ 4.5.2 Centrifugal-type Instability
Centrifugal instability is not restricted to the Görtler type; several curvature-driven variants may occur. In laminar separation bubbles, Gallaire2007 identified centrifugal mechanisms, and more recently, SavarinoJFM2025 reported similar effects in a transitional separation bubble. Such instabilities are modal in nature and have been associated with stationary eigenmodes (Gallaire2007; SavarinoJFM2025). Since the present low-frequency breathing originates in an eigenmode, a centrifugal mechanism represents a plausible explanation. To assess its likelihood, we evaluate the Rayleigh discriminant following the approach of Bayly1988, Sipp_Jacquin_PoF_2000, Gallaire2007, and SavarinoJFM2025.
@@ 5 Conclusion
Appendix A
This appendix describes the procedure used to determine the scaling of the eddy-viscosity field employed in the global stability and resolvent analyses.
The eddy viscosity is prescribed as
$\nu_{t}=\gamma,\nu_{t,\text{RANS}}(x,y),$ (17)
where $\nu_{t,\text{RANS}}$ denotes the eddy viscosity obtained from the RANS solution, and $\gamma$ is a scaling factor.
Using the unscaled RANS eddy viscosity ($\gamma=1$) yields unstable eigenvalues. Performing a resolvent analysis on this mean flow would therefore require the discounted resolvent formulation (Rolandi2024_invitation), which we aimed to avoid. A parametric study with different scaling factors was conducted to identify a suitable value of $\gamma$. Figure 19(a) shows the stationary eigenvalues ($\omega_{r}=0$) obtained for varying $\gamma$. Increasing $\gamma$ reduces the growth rates of the stationary eigenmodes; for $\gamma\geq 2$, all eigenvalues are stable.
The choice of $\gamma$ can be further guided by considering its influence on the frequency response. Bugeat2022JfM and SarrasPhD_thesis demonstrate that the growth rate of a stationary eigenmode determines the cutoff frequency of the associated low-pass filter in the resolvent response. This trend is evident in Fig. 19(b), which shows the resolvent gain for $\beta L_{b}/2\pi=1/3$ ($L/\lambda_{z}=1$), corresponding to the most energetic structure in the SPOD. As $\gamma$ increases beyond 2, the low-frequency plateau of the gain curve broadens, and the roll-off shifts to higher frequencies. This behaviour is characteristic of an increased low-pass filter cutoff frequency.
For the present analysis, $\gamma=3$ was selected as an appropriate compromise, providing a realistic cutoff frequency of $St\approx 0.01$ while avoiding the numerical difficulties associated with excessively high eddy-viscosity values.
Note that a more rigorous approach to avoid the tuning of eddy viscosity would involve assimilating the RANS solution to the measured mean field and incorporating a linearized eddy-viscosity model into the linear operatorSarrasJfM2024. This, however, is beyond the scope of the present study. The close agreement between SPOD and linearized results a posteriori justifies our approach.
Figure 19: (a) Stationary eigenvalues ($St=0$) computed for five eddy-viscosity scaling factors, $\nu_{t}(x,y)=\gamma,\nu_{t,\text{RANS}}(x,y)$ with $\gamma\in[1,2,3,4,5]$. (b) Leading resolvent gains for $\beta L_{b}/2\pi=1/3$ (equivalently $L/\lambda_{z}=1$) compared with the SPOD spectrum in the symmetry plane for the same set of $\gamma$ values.

### v3

@@ 4.5.2 Centrifugal-type Instability
Centrifugal instability is not restricted to the Görtler type; several curvature-driven variants may occur. In laminar separation bubbles, Barkley2002; Rodrguez2013 and Gallaire2007 identified centrifugal mechanisms, and more recently, SavarinoJFM2025 reported similar effects in a transitional separation bubble. Such instabilities are modal in nature and have been associated with stationary eigenmodes (Gallaire2007; Rodrguez2013; SavarinoJFM2025). Since the present low-frequency breathing originates in an eigenmode, a centrifugal mechanism represents a plausible explanation. To assess its likelihood, we evaluate the Rayleigh discriminant following the approach of Bayly1988, Sipp_Jacquin_PoF_2000, Barkley2002, Gallaire2007, and SavarinoJFM2025.
@@ Acknowledgments
Data availability statement.
The data that support the findings of this study are available from the corresponding author, LMF, upon reasonable request. The code used for the linearized analyses is an open source software package that is available at https://gitlab.com/felics-group/FELiCS.
Use of artificial intelligence (AI) tools.
Artificial intelligence tools (DeepL, ChatGPT, Perplexity) were used solely for language refinement.